# Swarm-based gradient descent meets simulated annealing

### v1

@@ 5.1. Convergence to the mean-field limit
$=$ $\mathbb{E}\left(\left|\frac{1}{N}\sum^{N}_{j=1}\left(m^{j}_{t}F(\mathbf{x}^{j}_{t})-\widetilde{m}^{j}_{t}F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)\right)\right|\right)$
$=$ $\mathbb{E}\left(\left|\frac{1}{N}\sum^{N}_{j=1}\left(m^{j}_{t}-\widetilde{m}^{j}_{t}\right)F(\mathbf{x}^{j}_{t})-\frac{1}{N}\sum^{N}_{j=1}\widetilde{m}^{j}_{t}\left(F(\mathbf{x}^{j}_{t})-F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)\right)\right|\right)$
$\leqslant$ $\mathbb{E}\sqrt{\frac{C}{N}\sum^{N}_{j=1}\left[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}+\left|m^{j}_{t}-\widetilde{m}^{j}_{t}\right|^{2}\right]}\,,$
where in the last inequality, we used the triangle inequality, Cauchy-Schwarz inequality, Lipschitz continuity of $F$ and upper boundedness of both $m$ [^4] Firstly we notice that, according to the Boundedness in $m$ condition in Assumption 2.1, the initial data for $m^{j}_{t}$ is bounded by $M_{\max}$, meaning $|m^{j}_{0}|<M_{\max}$ for all $j$ at the initial time. Given that $|F|\leqslant D$, then according to the $m^{j}_{t}$ equation in (1.2), we call Grönwall inequality to obtain that (5.5) $|m^{j}_{t}|\leqslant M_{\max}e^{2Dt}$ for all time. The same argument applies to bound $\widetilde{m}^{j}_{t}$. We denote the upper bound $c$ that depends on $M_{\max}$, $D$ and $t$. and $F$. The constant $C$ depends on $D$, $L$ and $M_{\max}$ (from Assumption 2.1).
Then controlling (5.3) and (5.4) gets translated to controlling the following term:
(5.6) $G(t)=\mathbb{E}\left(\frac{1}{N}\sum^{N}_{j=1}\left[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}+\left|m^{j}_{t}-\widetilde{m}^{j}_{t}\right|^{2}\right]\right)\,.$
If we can show $G(t)\sim 1/N$, the lemma is complete. The rest of the proof is dedicated to bounding $G(t)$. To do so, we will trace the evolution of $|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}$ and $|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}$ respectively.
  • To trace $|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}$, we compare the $\mathbf{x}^{j}_{t}$-equation in (1.2) and the $\widetilde{\mathbf{x}}^{j}_{t}$-equation in (5.1) to have:
@@
    (5.7) $\frac{\textnormal{d}}{\textnormal{d}t}\mathbb{E}[g(\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t})]$ $=\frac{\textnormal{d}}{\textnormal{d}t}\mathbb{E}[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}]\leqslant 2(L+L^{2})\left(\mathbb{E}[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}]+d\mathbb{E}[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}]\right)\,,$
    where we use Assumption 2.1 that both $\nabla F$ and $\sqrt{\sigma}$ are $L$-Lipschitz.
  • To trace $|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}$, we compare the $m$-equation in (1.2) and the $m$-equation in (5.1) to have:
    (5.8) $\frac{\textnormal{d}}{\textnormal{d}t}|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}$ $=-2(m^{j}_{t}-\widetilde{m}^{j}_{t})(m^{j}_{t}F(\mathbf{x}^{j}_{t})-\widetilde{m}^{j}_{t}F\left(\widetilde{\mathbf{x}}^{j}_{t}\right))$
    $\quad+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}m^{i}_{t}F(\mathbf{x}^{i}_{t})-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right)$
    $=-2(m^{j}_{t}-\widetilde{m}^{j}_{t})^{2}F(\mathbf{x}^{j}_{t})-2(m^{j}_{t}-\widetilde{m}^{j}_{t})\widetilde{m}^{j}_{t}(F(\mathbf{x}^{j}_{t})-F\left(\widetilde{\mathbf{x}}^{j}_{t}\right))$
    $\quad+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}m^{i}_{t}F(\mathbf{x}^{i}_{t})-\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})\right)$
    $\quad+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right)$
@@
    $\quad\underbrace{+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}(m^{i}_{t}-\widetilde{m}^{i}_{t})F(\mathbf{x}^{i}_{t})\right)}_{\text{(iii)}}$
    $\quad+\underbrace{2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}(F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t}))\right)}_{\text{(iv)}}$
    $\quad\underbrace{+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right)}_{\text{(v)}}\,.$
@@
      where $c$ comes from the boundedness of $m$, seen in Footnote 4 and $L$ is the Lipschitz constant from Assumption 2.1.
@@
      Similarly:
      $\begin{split}2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}(m^{i}_{t}-\widetilde{m}^{i}_{t})F(\mathbf{x}^{i}_{t})\right)&\leqslant|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\left(\frac{1}{N}\sum_{i=1}^{N}(m^{i}_{t}-\widetilde{m}^{i}_{t})F(\mathbf{x}^{i}_{t})\right)^{2}\\ &\leqslant D^{2}\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\left(\frac{1}{N}\sum_{i=1}^{N}(m^{i}_{t}-\widetilde{m}^{i}_{t})\right)^{2}\right]\\ &\leqslant D^{2}\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\frac{1}{N}\sum_{i=1}^{N}|m^{i}_{t}-\widetilde{m}^{i}_{t}|^{2}\right]\,,\end{split}$
@@
      $\begin{split}2(m^{j}_{t}-\widetilde{m}^{j}_{t})\Big{(}\frac{1}{N}\sum_{i=1}^{N}&\widetilde{m}^{i}_{t}(F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t}))\Big{)}\\ \leqslant&|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\Big{(}\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}(F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t}))\Big{)}^{2}\\ \leqslant&c\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\Big{(}\frac{1}{N}\sum_{i=1}^{N}(F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t}))\Big{)}^{2}\right]\\ \leqslant&c\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\frac{1}{N}\sum_{i=1}^{N}|F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t})|^{2}\right]\\ \leqslant&c\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\frac{L^{2}}{N}\sum_{i=1}^{N}|\mathbf{x}^{i}_{t}-\widetilde{\mathbf{x}}^{i}_{t}|^{2}\right]\end{split}$
      where we used Young’s in the first inequality, boundedness of $\widetilde{m}^{j}_{t}$ in the second (see Footnote 4), Hölder’s in the third, and finally the Lipschitz continuity of $F$ in the last.
@@
      $2(m^{j}_{t}-\widetilde{m}^{j}_{t})$ $\left(\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right)$
      $\leqslant$ $|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\left|\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right|^{2}\,.$
@@
    (5.10) $\frac{\textnormal{d}}{\textnormal{d}t}\left(\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}\right]\right)\leqslant\widetilde{C}\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}\left[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}+|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}\right]$
    $\hskip 144.54pt+\mathbb{E}\left[\left|\frac{1}{N}\sum_{j=1}^{N}\widetilde{m}^{j}_{t}F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right|^{2}\right]$
@@
Recall the definition of $G$ in (5.6), we collect (5.7) and (5.10) to have:
(5.11) $\frac{\textnormal{d}}{\textnormal{d}t}G(t)$ $\leqslant{C}(t)\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{N}|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}+\frac{1}{N}\sum_{j=1}^{N}|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}\right]$
$\qquad+\underbrace{\mathbb{E}\left[\left|\frac{1}{N}\sum_{j=1}^{N}\widetilde{m}^{j}_{t}F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right|^{2}\right]}_{\text{(vi)}}$
@@
where ${C}(t)$ depends on $D$, $M_{\max}$, $t$, $L$ and $d$. Noting that term (vi) can be further controlled using the law of large numbers: let $X=mF(\mathbf{x})$ be a random variable, then $\widetilde{m}^{j}_{t}F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)$ is a realization of $X$ and we have
$\overline{X}_{N}=\frac{1}{N}\sum_{j=1}^{N}\widetilde{m}^{j}_{t}F(\widetilde{\mathbf{x}}^{j}_{t})$
then, since $\{(\widetilde{\mathbf{x}}^{j}_{t},\widetilde{m}^{j}_{t})\}_{j=1}^{N}$ are i.i.d. samples of $\mu$ we have
[…]
Combining with (5.11), we have:
$\frac{\textnormal{d}}{\textnormal{d}t}G(t)\leqslant C(t)G(t)+\mathcal{O}\left(\frac{1}{N}\right)\quad\text{with}\quad G(0)=0\,,$
then, calling Grönwall’s inequality, we have $G(t)\leqslant\mathcal{O}\left(\frac{1}{N}\right),$ which completes the proof. ∎

### v2

@@ 5.1. Convergence to the mean-field limit
$\leqslant$ $\mathbb{E}\left(|\widetilde{F}^{N}_{t}-\overline{F}^{N}_{t}|\big{|}\Omega\right)+\frac{C}{N}$
$\leqslant$ $C\underbrace{\mathbb{E}\left(\sqrt{\frac{1}{N}\sum^{N}_{j=1}\left[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}+\left|m^{j}_{t}-\widetilde{m}^{j}_{t}\right|^{2}\right]}\middle|\Omega\right)}_{\leqslant\sqrt{G(t)}}+C\underbrace{\mathbb{E}\left(\left|\frac{1}{N}\sum m^{j}_{0}-\mathbb{E}_{\mu_{0}}[m]\right|\middle|\Omega\right)}_{\leqslant\frac{1}{\mathbb{P}(\Omega)}\mathbb{E}\left(\left|\frac{1}{N}\sum m^{j}_{0}-\mathbb{E}_{\mu_{0}}[m]\right|\right)\leqslant\frac{C}{\sqrt{N}}}+\frac{C}{N}\,.$
Once again, if we can show $G(t)\sim 1/N$, (5.4) is complete.
The rest of the proof is dedicated to showing the bound for $G(t)$. To do so, we will trace the evolution of $|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}$ and $|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}$ respectively.
  • To trace $|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}$, we compare the $\mathbf{x}^{j}_{t}$-equation in (SSA) and the $\widetilde{\mathbf{x}}^{j}_{t}$-equation in (5.1) to have:
@@
    (5.10) $\frac{\textnormal{d}}{\textnormal{d}t}\mathbb{E}[g(\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t})|\Omega]$ $=\frac{\textnormal{d}}{\textnormal{d}t}\mathbb{E}[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}|\Omega]\leqslant 2(L+L^{2})\left(\mathbb{E}[|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}|\Omega]+d\mathbb{E}[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}|\Omega]\right)\,,$
    where we used Assumption 2.1 that both $\nabla F$ and $\sqrt{\sigma}$ are $L$-Lipschitz.
  • To trace $|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}$, we compare the $m$-equation in (SSA) and the $m$-equation in (5.1) to have:
    (5.11) $\frac{\textnormal{d}}{\textnormal{d}t}|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}$ $=-2(m^{j}_{t}-\widetilde{m}^{j}_{t})(m^{j}_{t}F(\mathbf{x}^{j}_{t})-\widetilde{m}^{j}_{t}F(\widetilde{\mathbf{x}}^{j}_{t}))$
    $\quad+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{\frac{1}{N}\sum_{i=1}^{N}m^{i}_{t}F(\mathbf{x}^{i}_{t})}{\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}}-\frac{\mathbb{E}_{\mu}[mF(\mathbf{x})]}{\mathbb{E}_{\mu}(m)}\right)$
    $=-2(m^{j}_{t}-\widetilde{m}^{j}_{t})^{2}F(\mathbf{x}^{j}_{t})-2(m^{j}_{t}-\widetilde{m}^{j}_{t})\widetilde{m}^{j}_{t}\left(F(\mathbf{x}^{j}_{t})-F(\widetilde{\mathbf{x}}^{j}_{t})\right)$
    $\quad+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}m^{i}_{t}F(\mathbf{x}^{i}_{t})-\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})\right)/\left(\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}\right)$
    $\quad+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})}{\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}}-\frac{\mathbb{E}_{\mu}[mF(\mathbf{x})]}{\mathbb{E}_{\mu}(m)}\right)\,.$
@@
    $\quad\underbrace{+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}(m^{i}_{t}-\widetilde{m}^{i}_{t})F(\mathbf{x}^{i}_{t})\right)/\left(\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}\right)}_{\text{(iii)}}$
    $\quad+\underbrace{2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}(F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t}))\right)/\left(\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}\right)}_{\text{(iv)}}$
    $\quad\underbrace{+2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})}{\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}}-\frac{\mathbb{E}_{\mu}[mF(\mathbf{x})]}{\mathbb{E}_{\mu}(m)}\right)}_{\text{(v)}}\,.$
@@
      where $c$ comes from the boundedness of $m$, seen in Footnote 5 and $L$ is the Lipschitz constant from Assumption 2.1.
@@
      Because of the total mass preserving and the definition of $\Omega$, we have
      $\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}=\frac{1}{N}\sum^{N}_{i=1}m^{i}_{0}\geqslant\frac{1}{2}\,.$
      Thus,
      $\begin{split}&2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}(m^{i}_{t}-\widetilde{m}^{i}_{t})F(\mathbf{x}^{i}_{t})\right)/\left(\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}\right)\\ \leqslant&2|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+2\left(\frac{1}{N}\sum_{i=1}^{N}(m^{i}_{t}-\widetilde{m}^{i}_{t})F(\mathbf{x}^{i}_{t})\right)^{2}\\ \leqslant&D^{2}\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\left(\frac{1}{N}\sum_{i=1}^{N}(m^{i}_{t}-\widetilde{m}^{i}_{t})\right)^{2}\right]\\ \leqslant&D^{2}\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\frac{1}{N}\sum_{i=1}^{N}|m^{i}_{t}-\widetilde{m}^{i}_{t}|^{2}\right]\,,\end{split}$
@@
      $\begin{split}&2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}(F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t}))\right)/\left(\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}\right)\\ \leqslant&2|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+2\left(\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}(F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t}))\right)^{2}\\ \leqslant&c\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\left(\frac{1}{N}\sum_{i=1}^{N}(F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t}))\right)^{2}\right]\\ \leqslant&c\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\frac{1}{N}\sum_{i=1}^{N}|F(\mathbf{x}^{i}_{t})-F(\widetilde{\mathbf{x}}^{i}_{t})|^{2}\right]\\ \leqslant&c\left[|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+\frac{L^{2}}{N}\sum_{i=1}^{N}|\mathbf{x}^{i}_{t}-\widetilde{\mathbf{x}}^{i}_{t}|^{2}\right]\end{split}$
      where we used Young’s inequality in the first inequality, boundedness of $\widetilde{m}^{j}_{t}$ in the second (see Footnote 5), Hölder’s in the third, and finally the Lipschitz continuity of $F$ in the last inequality.
@@
      $2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})}{\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}}-\frac{\mathbb{E}_{\mu}[mF(\mathbf{x})]}{\mathbb{E}_{\mu}(m)}\right)$
      $=$ $2(m^{j}_{t}-\widetilde{m}^{j}_{t})\left(\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right)/\left(\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}\right)$
      $+2(m^{j}_{t}-\widetilde{m}^{i}_{t})\mathbb{E}_{\mu}[mF(\mathbf{x})]\left(\frac{1}{\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}}-\frac{1}{\mathbb{E}_{\mu}(m)}\right)$
      $\leqslant$ $4|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}+2\left|\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right|^{2}+2D^{2}\left|\frac{1}{N}\sum^{N}_{i=1}m^{i}_{0}-\mathbb{E}_{\mu_{0}}(m)\right|^{2}\,,$
      where we use $\frac{1}{N}\sum^{N}_{i=1}m^{i}_{t}=\frac{1}{N}\sum^{N}_{i=1}m^{i}_{0}\geqslant\frac{1}{2}$, $\mathbb{E}_{\mu}(m)=\mathbb{E}_{\mu_{0}}(m)$, and boundedness of $F$ in the last inequality.
@@
    (5.13) $\frac{\textnormal{d}}{\textnormal{d}t}\left(\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{N}|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}\middle|\Omega\right]\right)$
    $\leqslant\widetilde{C}\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{N}|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}+\frac{1}{N}\sum_{j=1}^{N}|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}\middle|\Omega\right]$
    $\quad+\widetilde{C}\mathbb{E}\left[\left|\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right|^{2}+\left|\frac{1}{N}\sum^{N}_{i=1}m^{i}_{0}-\mathbb{E}_{\mu_{0}}(m)\right|^{2}\middle|\Omega\right]\,,$
@@
Recall the definition of $G$ in (5.6). We collect (5.10) and (5.13) to have:
(5.14) $\frac{\textnormal{d}}{\textnormal{d}t}G(t)$ $\leqslant C(t)\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{N}|\mathbf{x}^{j}_{t}-\widetilde{\mathbf{x}}^{j}_{t}|^{2}+\frac{1}{N}\sum_{j=1}^{N}|m^{j}_{t}-\widetilde{m}^{j}_{t}|^{2}\middle|\Omega\right]$
$\qquad+C(t)\underbrace{\mathbb{E}\left[\left|\frac{1}{N}\sum_{i=1}^{N}\widetilde{m}^{i}_{t}F(\widetilde{\mathbf{x}}^{i}_{t})-\mathbb{E}_{\mu}[mF(\mathbf{x})]\right|^{2}+\left|\frac{1}{N}\sum^{N}_{i=1}m^{i}_{0}-\mathbb{E}_{\mu_{0}}(m)\right|^{2}\middle|\Omega\right]}_{\text{(vi)}}$
@@
where ${C}(t)$ depends on $D$, $M_{\max}$, $t$, $L$ and $d$. Noting that term (vi) can be further controlled using the law of large numbers: for the first term, let $X=mF(\mathbf{x})$ be a random variable. Then, $\widetilde{m}^{j}_{t}F\left(\widetilde{\mathbf{x}}^{j}_{t}\right)$ is a realization of $X$ and we have
$\overline{X}_{N}=\frac{1}{N}\sum_{j=1}^{N}\widetilde{m}^{j}_{t}F(\widetilde{\mathbf{x}}^{j}_{t})\,.$
Since $\{(\widetilde{\mathbf{x}}^{j}_{t},\widetilde{m}^{j}_{t})\}_{j=1}^{N}$ are i.i.d. samples of $\mu$, we have
$\mathbb{E}[|\overline{X}_{N}-\mathbb{E}[X]|^{2}|\Omega]\leqslant\mathbb{E}[|\overline{X}_{N}-\mathbb{E}[X]|^{2}]/\mathbb{P}(\Omega)\leqslant\text{Var}[\overline{X}_{N}]/\mathbb{P}(\Omega)\leqslant\frac{1}{N}\text{Var}[X]/\mathbb{P}(\Omega)=\mathcal{O}\left(\frac{1}{N}\right)\,,$
where we use $\mathbb{P}(\Omega)\geqslant 1-\frac{C}{N}$. Similar bound also holds for the second term in (vi). Combining with (5.14), we have:
$\frac{\textnormal{d}}{\textnormal{d}t}G(t)\leqslant C(t)G(t)+\mathcal{O}\left(\frac{1}{N}\right)\quad\text{with}\quad G(0)=0\,.$
Finally, calling Grönwall’s inequality, we have $G(t)\leqslant\mathcal{O}\left(\frac{1}{N}\right),$ which completes the proof. ∎